### v1

# On the Fujita Phenomenon for a Forced Spatio-Temporal Fractional Diffusion Equation

Abstract.
We study the Cauchy problem for a semilinear fractional diffusion equation with a time-dependent forcing term:
$\partial_{t}^{\alpha}u+(-\Delta)^{\mathsf{s}}u=|u|^{p}+t^{\sigma}\,\mathbf{w}(x),\quad(t,x)\in(0,\infty)\times\mathbb{R}^{N},$
with parameters $\alpha,\mathsf{s}\in(0,1)$, $\sigma>-\alpha$, and a continuous function $\mathbf{w}$. The operator $\partial_{t}^{\alpha}$ denotes the Caputo fractional derivative. Our main contributions are threefold. First, we prove the local existence of mild solutions and demonstrate a finite-time blow-up in the subcritical regime, provided $\int_{\mathbb{R}^{N}}\mathbf{w}(x)\,dx>0$. Second, for the supercritical case with $\alpha<\sigma<0$, we establish the global existence for sufficiently small initial data and a forcing term, and we determine the critical exponent to be
$p_{F}=\frac{N\alpha-2\mathsf{s}\sigma}{N\alpha-2\mathsf{s}(\alpha+\sigma)}.$
Finally, for this supercritical range, we prove a more robust global existence result under assumptions requiring only local smallness and controlled growth of the data.
Key words and phrases:
Fractional diffusion equation, local/global existence, blow-up, Fujita exponent.
2020 Mathematics Subject Classification:
35K58, 35R11, 35K15, 35C15, 35B44, 35B33, 35A01, 35B60
1. Introduction and main results
In this paper, we study the Cauchy problem for a spatio-temporal fractional diffusion equation with a forcing term:
$\begin{cases}\partial_{t}^{\alpha}u+(-\Delta)^{\mathsf{s}}u=|u|^{p}+t^{\sigma}\mathbf{w}(x)&\text{in }\mathbb{R}^{N}\times(0,\infty),\\ u(x,0)=u_{0}(x)&\text{in }\mathbb{R}^{N},\end{cases}$ (1.1)
where $\alpha,\mathsf{s}\in(0,1)$, $\sigma>-1$, $p>1$, and $\mathbf{w}$ is a given continuous function. In what follows, $\partial_{t}^{\alpha}$ for $\alpha\in(0,1)$ is the Caputo fractional derivative,
$\partial_{t}^{\alpha}u(t,x)=\frac{1}{\Gamma(1-\alpha)}\int_{0}^{t}(t-\tau)^{-\alpha}\partial_{\tau}u(\tau,x)\,d\tau,$
and $(-\Delta)^{\mathsf{s}}$ for $\mathsf{s}\in(0,1)$ is the fractional Laplacian, defined spectrally as $(-\Delta)^{\mathsf{s}}u:=\mathcal{F}^{-1}\big(|\xi|^{2\mathsf{s}}\mathcal{F}(u)\big)$ with $\mathcal{F}$ denoting the Fourier transform [1, 30, 37, 38].
Equations that are fully nonlocal, featuring both fractional time derivatives and fractional spatial operators, provide an effective framework for modeling phenomena with long-range interactions and memory effects. Such models naturally occur in a variety of physical settings, including plasma physics and turbulent transport—where anomalous diffusion arises from long-range correlations in space and time [9, 10]—geophysical flows and atmospheric dynamics, where memory and large-scale interactions play a fundamental role [34], as well as in financial mathematics for asset price dynamics incorporating long-term memory and jump processes [8].
From the analytical perspective, a central theme is the study of the critical exponent that delineates global existence from finite-time blow-up. The classical starting point is the seminal work of Fujita [17, 18] on the semilinear heat equation (corresponding to $\alpha=1$, $\mathsf{s}=1$ and $\mathbf{w}\equiv 0$ in our setting). Fujita proved that for $1<p<1+\frac{2}{N}$ and nontrivial initial data $u_{0}\not\equiv 0$, every solution blows up in finite time, while for $p>1+\frac{2}{N}$ global solutions exist for sufficiently small initial data. The critical value $p_{F}=1+\frac{2}{N}$ was later also shown to lie in the blow-up regime [22, 27]. These foundational results are thoroughly reviewed in Levine’s survey [13].
The Fujita phenomenon has been widely extended beyond the classical setting. Zhang and Sun [39] studied Problem (1.1) without a forcing term ($\mathbf{w}=0$) in the case $\mathsf{s}=1$ and $0<\alpha<1$, and established an analogous critical exponent $p_{F}=1+\frac{2}{N}$. For the fully fractional situation, Kirane et al. [21] analyzed the more general equation
$\partial_{t}^{\alpha}u+(-\Delta)^{\mathsf{s}}u=h(x,t)\,|u|^{p},$ (1.2)
under the assumption $h(x,t)\geq C_{h}|x|^{\sigma}t^{\rho}$. They showed that no global solutions exist whenever
$1<p\leq 1+\frac{\alpha(2\mathsf{s}+\sigma)+2\mathsf{s}\rho}{\alpha N+2\mathsf{s}(1-\alpha)},$
thereby extending the Fujita-type blow-up mechanism to equations involving both time-fractional and space-fractional operators.
This direction of research fits within a broad program aimed at characterizing critical exponents for nonlinear evolution equations. In particular, Guedda and Kirane [19, 20] developed a unified methodology for proving nonexistence of global solutions across a wide class of problems. Their framework includes the case $\alpha=1$ in (1.2), and clarifies how the interplay between the coefficient $h$, the parameters $p$, $\mathsf{s}$, and the spatial dimension $N$ determines the critical threshold. Their approach, relying on carefully chosen scaled test functions, yields a powerful and flexible tool for identifying critical exponents influenced by both the diffusion operator and the growth of the nonlinearity.
For the unforced version of our main problem (1.1) (that is, $\mathbf{w}=0$), the critical exponent can be inferred from the scaling properties of the equation. If $u(t,x)$ is a solution, then for any $\lambda>0$ the rescaled function
$u_{\lambda}(t,x):=\lambda^{\frac{2\mathsf{s}}{p-1}}u\!\left(\lambda^{\frac{2\mathsf{s}}{\alpha}}t,\,\lambda x\right)$
is again a solution. The corresponding scale-invariant Lebesgue space for the initial data is $L^{p_{c}}(\mathbb{R}^{N})$, where
$p_{c}=\frac{N(p-1)}{2\mathsf{s}}.$ (1.3)
This leads to the Fujita-type critical exponent
$p_{F}:=1+\frac{2\mathsf{s}}{N},$
which reduces to the classical value when $\alpha=\mathsf{s}=1$; see [16].
It is, however, well understood that critical exponents are highly sensitive to the structure of the problem. The addition of a forcing term may substantially alter the threshold between global existence and blow-up. For the classical heat equation, for example, forcing terms can shift the Fujita exponent [5]. Likewise, geometric effects may change the critical value, as demonstrated by Bandle and Levine [4] for conical domains.
Motivated by these considerations, and drawing on the general criticality framework developed in [19], we examine how the forcing term $t^{\sigma}\mathbf{w}(x)$ affects the Fujita exponent in the fully fractional setting of equation (1.1). Our objective is to determine whether, and in what manner, the critical exponent
$p_{F}=1+\frac{2\alpha}{N\alpha+2(1-\alpha)}$
obtained for the homogeneous case is modified in the presence of an external source. Our main result establishes a sharp critical exponent for the forced equation, precisely characterizing the transition between global existence and finite-time blow-up.
We begin by establishing local well-posedness in the Lebesgue space $L^{q}$ under appropriate assumptions.
Theorem 1.1.
Let $p>1$, $\alpha,\mathsf{s}\in(0,1)$, $\sigma>-\alpha$, and $q\in(p_{c},\infty]$ with $q\geq p$, where the critical exponent $p_{c}$ is given by (1.3). Assume that the initial data $u_{0}$ and the forcing term $\mathbf{w}$ both belong to $L^{q}(\mathbb{R}^{N})$. Then, there exists a time $T>0$ such that Problem (1.1) admits a unique solution
$u\in L^{\infty}(0,T;L^{q}(\mathbb{R}^{N})).$
Remark 1.1.
  1. (i)
    Local existence may also be obtained via semigroup methods. More precisely, one can introduce the solution operators $\mathbb{P}_{\alpha,\mathsf{s}}(t)$ and $\mathbb{S}_{\alpha,\mathsf{s}}(t)$ through the subordination principle, following [3]:
    $\mathbb{P}_{\alpha,\mathsf{s}}(t)u_{0}$ $=\int_{0}^{\infty}\phi_{\alpha}(\theta)\,T_{\mathsf{s}}(t^{\alpha}\theta)\,u_{0}\,d\theta,$
    $\mathbb{S}_{\alpha,\mathsf{s}}(t)\varphi$ $=\alpha\int_{0}^{\infty}\theta\,\phi_{\alpha}(\theta)\,T_{\mathsf{s}}(t^{\alpha}\theta)\,\varphi\,d\theta,$
    where $T_{\mathsf{s}}(t)=e^{-t(-\Delta)^{\mathsf{s}}}$ denotes the fractional heat semigroup and $\phi_{\alpha}$ is the Wright-type function (see [31]). This approach yields mild solutions in $C([0,T],L^{q}(\mathbb{R}^{N}))$. In the present work, however, we rely on the Duhamel representation involving the kernels $Z$ and $Y$, which is better adapted to capturing spatial decay and regularity effects.
  2. (ii)
    In the case $\mathsf{s}=1$ and $\mathbf{w}=0$, Zhang and Sun [39] established local well-posedness for initial data in $C_{0}(\mathbb{R}^{N})$ and provided a complete characterization of the critical exponent governing the transition between finite-time blow-up and global existence.
  3. (iii)
    The classical situation $\alpha=\mathsf{s}=1$ was previously studied in [23] under the assumptions $\sigma>-1$ and $\sigma\neq 0$. In that setting, both local well-posedness and the corresponding critical exponent were rigorously derived.
Our next result addresses the nonexistence of global solutions to (1.1). The precise statement is as follows:
Theorem 1.2.
Let $N\geqslant 1$ and $\alpha,\mathsf{s}\in(0,1)$. Assume further that ${\mathbf{w}}\in C_{0}(\mathbb{R}^{N})\cap L^{1}(\mathbb{R}^{N})$ satisfies
$\int\limits_{\mathbb{R}^{N}}{\mathbf{w}}(x)\,dx>0.$ (1.4)
  1. (i)
    If $\alpha<\sigma\leq 0$, $2\mathsf{s}<N$, and
    $1<p<p^{*}\vcentcolon=\frac{N\alpha-2\mathsf{s}\sigma}{N\alpha-2\mathsf{s}(\alpha+\sigma)},$ (1.5)
    then problem (1.1) admits no global weak solution.
  2. (ii)
    If $\sigma>0$, then the same conclusion holds for every $p>1$.
Remark 1.2.
  1. (i)
    The special case $\alpha=1$ with $\sigma=0$ was recently analyzed in [29]. A broader framework for the case $\alpha=1$ is developed in [7].
  2. (ii)
    For additional nonlocal settings, we refer the reader to [6].
  3. (iii)
    The condition $N>2\mathsf{s}$ ensures that the exponent $p^{*}$ defined in (1.5) satisfies $p^{*}>1$.
  4. (iv)
    For related contributions in the setting $\alpha=\mathsf{s}=1$, we refer the reader to [2, 5, 23, 33]. In particular, when $\sigma\in(-1,0)$, one has
    $p^{*}=\frac{N-2\sigma}{\,N-2\sigma-2\,}$,
    which is consistent with [23, Theorem 1.1, (1.8)].
  5. (v)
    The special case $\alpha=1$ with integer $\mathsf{s}\geq 1$ was previously analyzed in [32].
  6. (vi)
    Our results show that the presence of the external forcing term $t^{\sigma}\mathbf{w}(x)$ accelerates finite-time blow-up compared to the dynamics observed in the unforced case.
  7. (vii)
    More specifically, the analysis indicates that the space-time dependent forcing term counteracts the regularizing effects of the spatio-temporal diffusion governed by the fractional operators $\partial_{t}^{\alpha}$ and $(-\Delta)^{\mathsf{s}}$.
  8. (viii)
    A crucial factor for blow-up is the positivity of the total mass of $\mathbf{w}$. In particular, when $\sigma>0$, the temporal amplification enhances the effect of the forcing, thereby lowering the critical exponent $p^{*}$ beyond which solutions necessarily blow up.
  9. (ix)
    These results emphasize the delicate balance between dissipation due to fractional diffusion and amplification induced by external forcing, highlighting the significant role of the temporal factor $t^{\sigma}$ in governing blow-up.
We now turn to the analysis of the global theory. Global existence is established under the assumption that the initial data is sufficiently small in the critical norm, and that the forcing term is suitably small in an appropriate functional norm. The main result in this setting is stated as follows:
Theorem 1.3.
Let $N\geq 1$, $\alpha,\mathsf{s}\in(0,1)$, and $\alpha<\sigma\leq 0$. Assume that $p\geq\frac{N\alpha-2\mathsf{s}\sigma}{N\alpha-2\mathsf{s}(\alpha+\sigma)}$, where $N>2\mathsf{s}$. Let us define ${r_{c}}=\frac{N\alpha(p-1)}{2\mathsf{s}\sigma(p-1)+2\mathsf{s}\alpha p}$ and recall that $p_{c}$ is given by (1.3). Then, for any initial data $u_{0}\in L^{p_{c}}(\mathbb{R}^{N})$ and $\mathbf{w}\in L^{r_{c}}(\mathbb{R}^{N})$, with the property that $\|u_{0}\|_{L^{p_{c}}}+\|\mathbf{w}\|_{L^{r_{c}}}$ is sufficiently small, the equation (1.1) admits a global solution $u$ in time.
Remark 1.3.
  1. (i)
    When $\sigma=0$, the scaling properties imply that $L^{p_{c}}(\mathbb{R}^{N})$ is the unique Lebesgue space with the following feature: if the initial data are small in this space and the forcing term $\mathbf{w}$ is small in $L^{\frac{p_{c}}{p}}(\mathbb{R}^{N})$, then a global solution exists in $L^{q}$.
  2. (ii)
    In the absence of a forcing term (i.e., $\mathbf{w}=0$), the problem is addressed in [11]. There, global existence is established either under a strong condition, requiring the smallness of $\|u_{0}\|_{L^{p_{c}}(\mathbb{R}^{N})}$, or under a weaker condition, relying on smallness of local averages of $u_{0}$.
  3. (iii)
    The global existence result can be extended to $L^{q}$ spaces outside the distinguished range, provided additional integrability conditions are imposed on $u_{0}$. In the critical case $p^{*}=1+\frac{2\mathsf{s}}{N}$, the existence of global solutions depends on $\alpha$: for $\alpha\in(0,1)$, global solutions exist, whereas for $\alpha=1$, every nonnegative nontrivial solution blows up in finite time.
When a forcing term depends on both time and space, a weaker criterion for the global existence of $L^{q}$-solutions can be formulated by imposing smallness conditions on the local averages of the initial data $u_{0}$ and the forcing term $\mathbf{w}$ over balls of suitable radius.
Theorem 1.4.
Let $\alpha,\mathsf{s}\in(0,1)$, $\alpha<\sigma<0$, and $2\mathsf{s}<N$. Assume
$\frac{N\alpha-2\mathsf{s}\sigma}{N\alpha-2\mathsf{s}(\alpha+\sigma)}\leq p<\frac{N\alpha-2\mathsf{s}q\sigma}{N\alpha-2\mathsf{s}q(\alpha+\sigma)}$
and let $q$ satisfy
$p_{c}<q\leq pp_{c},\quad q\geq p,$
where $p_{c}$ is given by (1.3). Let $u_{0},\mathbf{w}\in L^{p_{c}}(\mathbb{R}^{N})\cap L^{q}(\mathbb{R}^{N})$, with $u_{0},\mathbf{w}\geq 0$. Assume the scaling condition
$(\sigma+\alpha)<0\leq\beta:=\frac{\alpha}{p-1}-\frac{N\alpha}{2\mathsf{s}q}<\frac{1}{p}.$
Let $R_{0}=\min\left\{\left(\frac{\mathscr{M}}{2\|u_{0}\|_{L^{q}}}\right)^{\frac{1}{\beta}},\left(\frac{\mathscr{M}}{2C\|\mathbf{w}\|_{L^{q}}}\right)^{\frac{1}{\beta+\alpha+\sigma}}\right\}$ with $\mathscr{M}$ from Lemma 2.6, and choose $M\geq 0$, $\delta>0$ such that
$\|u_{0}\|_{L^{p_{c}}(|x|>MR_{0}^{\alpha/2\mathsf{s}})}\leq\frac{\mathscr{M}}{4\|F\|_{L^{r}}},$ (1.6)
$\|w\|_{L^{p_{c}}(|x|>MR_{0}^{\alpha/2\mathsf{s}})}\leq\frac{\mathscr{M}}{4\|F\|_{L^{r}}},$ (1.7)
$\|F\|_{L^{r}(|\xi|<\delta)}\leq\frac{\mathscr{M}}{\max\left(\|u_{0}\|_{L^{p_{c}}},\|\mathbf{w}\|_{L^{p_{c}}}\right)},$ (1.8)
with $1-\dfrac{1}{r}=\dfrac{1}{p_{c}}-\dfrac{1}{q}$, If the local averages satisfy
$\sup_{R>MR_{0}}R^{\frac{\alpha}{p-1}-\frac{N\alpha}{2\mathsf{s}}}\int\limits_{|y|<R^{\frac{\alpha}{2\mathsf{s}}}}u_{0}(y)\,dy<\frac{\mathscr{M}M^{\frac{2\mathsf{s}}{p-1}-N}}{8\|F\|_{L^{q}(|\xi|>\delta)}},$ (1.9)
$\sup_{R>MR_{0}}R^{\frac{\alpha}{p-1}-\frac{N\alpha}{2\mathsf{s}}+\sigma+\alpha}\int\limits_{|y|<R^{\frac{\alpha}{2\mathsf{s}}}}\mathbf{w}(y)\,dy<\frac{\mathscr{M}M^{\frac{2\mathsf{s}}{p-1}-N}}{8\|F\|_{L^{q}(|\xi|>\delta)}},$ (1.10)
then the maximal $L^{q}$-solution exists globally in time.
The two global existence results, Theorem 1.3 and Theorem 1.4, are established under distinct sets of assumptions. In Theorem 1.3, global existence is ensured by imposing smallness conditions on the norms of the initial data $u_{0}$ and the forcing term $\mathbf{w}$. Specifically, these smallness conditions are formulated in terms of norms associated with the critical exponent $p_{c}$ and the exponent $r_{c}$ introduced in the proof.
By contrast, Theorem 1.4 does not require smallness in any norm. Instead, it relies on assumptions concerning the behavior and decay of local averages of $u_{0}$ and $\mathbf{w}$ over balls. This approach emphasizes the spatial distribution and asymptotic properties of the data rather than their global magnitude.
Therefore, while Theorem 1.3 is grounded in quantitative smallness within specific functional spaces, Theorem 1.4 provides an alternative existence criterion based on qualitative spatial decay and distribution properties.
Outline of the paper. The paper is organized as follows. Section 2 collects preliminary material. We begin by reviewing essential concepts related to fractional derivatives and then introduce the fundamental kernels used to construct mild solutions via the Duhamel formula. The key properties of these kernels—self-similarity, integrability, and regularity—form the basis of our analysis of the existence and qualitative behavior of solutions. This section also establishes several technical lemmas that will serve as crucial tools for the subsequent estimates. In Section 3, we provide the proofs of the main results. We first establish a local existence result as stated in Theorem 1.1. Next, we address the nonexistence of global solutions by proving Theorem 1.2. We then prove the global existence result in Theorem 1.3, which relies on the smallness of the initial data $u_{0}$ in the critical norm and the smallness of the forcing term $\mathbf{w}$ in a suitable $L^{r_{c}}$ norm. Finally, we present Theorem 1.4, establishing global existence under significantly weaker local smallness conditions. Unlike Theorem 1.3, this result allows $u_{0}$ and $\mathbf{w}$ to be large in a global norm, provided they satisfy appropriate local smallness and decay assumptions.
Throughout the remainder of the article, the constant $C$ may denote different values in different contexts.
2. Useful tools & Auxiliary results
In this section, we introduce the notation used throughout the paper and present several auxiliary results and estimates.
We begin with a brief introduction to the Caputo fractional derivative. For a comprehensive treatment of fractional calculus, we refer the reader to [14, 25, 35, 36]. The Riemann–Liouville fractional integral of order $\alpha\in(0,1)$ is defined by
$\mathcal{I}^{\alpha}u(t)=\frac{1}{\Gamma(\alpha)}\int\limits_{0}^{t}(t-s)^{\alpha-1}u(s)\,ds,$ (2.1)
where $\Gamma$ denotes the Gamma function, given by
$\Gamma(z)=\int\limits_{0}^{\infty}s^{z-1}e^{-s}\,ds,\quad\text{for }\Re(z)>0.$
The Riemann–Liouville fractional derivative of order $\alpha\in(0,1)$ is then defined as
$\mathscr{D}^{\alpha}u(t)=\frac{d}{dt}\left(\mathcal{I}^{1-\alpha}u(t)\right)=\frac{1}{\Gamma(1-\alpha)}\frac{d}{dt}\left(\int\limits_{0}^{t}(t-s)^{-\alpha}u(s)\,ds\right),$ (2.2)
provided that $t\mapsto\mathcal{I}^{1-\alpha}u(t)$ is absolutely continuous.
The Caputo fractional derivative of order $\alpha\in(0,1)$ is defined by
$\partial_{t}^{\alpha}u(t)=\mathscr{D}^{\alpha}\left(u(t)-u(0)\right),$ (2.3)
where $u(0)$ exists and $t\mapsto\mathcal{I}^{1-\alpha}u(t)$ is absolutely continuous.
Under the stronger assumption that $u$ is absolutely continuous on $[0,T]$, the Caputo derivative admits the following equivalent representation (see [25, Theorem 2.1, p. 92]):
$\partial_{t}^{\alpha}u(t)=\frac{1}{\Gamma(1-\alpha)}\int\limits_{0}^{t}(t-s)^{-\alpha}\frac{du}{ds}(s)\,ds.$ (2.4)
We now state the following integration-by-parts formula holds for the Caputo derivative:
Lemma 2.1 (Integration by parts).
Let $\partial_{t}^{\alpha}u\in L^{1}(0,T)$, $v\in C([0,T])$, and $v(T)=0$. Then,
$\int\limits_{0}^{T}v(t)\partial_{t}^{\alpha}u(t),dt=\int\limits_{0}^{T}\left(u(t)-u(0)\right)\mathscr{D}_{t|T}^{\alpha}v(t)dt,$ (2.5)
where $\mathscr{D}_{t|T}^{\alpha}$ is the right-sided Riemann-Liouville fractional derivative of order $\alpha$, defined by
$\mathscr{D}_{t|T}^{\alpha}v(t)=\frac{1}{\Gamma(1-\alpha)}\frac{d}{dt}\int\limits_{t}^{T}(t-s)^{-\alpha}v(s)ds.$ (2.6)
Proof.
Using the definition of the Caputo derivative, we write
$\int\limits_{0}^{T}v(t)\,\partial_{t}^{\alpha}u(t)\,dt=\frac{1}{\Gamma(1-\alpha)}\int\limits_{0}^{T}v(t)\left(\int\limits_{0}^{t}(t-s)^{-\alpha}u^{\prime}(s)\,ds\right)dt.$
By Fubini’s theorem, we interchange the order of integration:
$\frac{1}{\Gamma(1-\alpha)}\int\limits_{0}^{T}u^{\prime}(s)\left(\int\limits_{s}^{T}(t-s)^{-\alpha}v(t)\,dt\right)ds=\int\limits_{0}^{T}u^{\prime}(s)\,I^{1-\alpha}_{s|T}v(s)\,ds,$
where $I^{1-\alpha}_{s|T}v(s)=\frac{1}{\Gamma(1-\alpha)}\int\limits_{s}^{T}(t-s)^{-\alpha}v(t)\,dt$.
We now integrate by parts, noting that $\frac{d}{ds}I^{1-\alpha}_{s|T}v(s)=-\mathscr{D}_{s|T}^{\alpha}v(s)$, and use the fact that $v(T)=0$:
$\int\limits_{0}^{T}u^{\prime}(s)\,I^{1-\alpha}_{s|T}v(s)\,ds=-\int\limits_{0}^{T}u(s)\,\mathscr{D}_{s|T}^{\alpha}v(s)\,ds+\left[u(s)I^{1-\alpha}_{s|T}v(s)\right]_{0}^{T}.$
Since $I^{1-\alpha}_{T|T}v(T)=0$, we obtain
$\int\limits_{0}^{T}v(t)\,\partial_{t}^{\alpha}u(t)\,dt=-\int\limits_{0}^{T}u(t)\,\mathscr{D}_{t|T}^{\alpha}v(t)\,dt+u(0)I^{1-\alpha}_{0|T}v(0).$
It remains to compute the integral of the right-sided Riemann–Liouville derivative:
$\int\limits_{0}^{T}\mathscr{D}_{t|T}^{\alpha}v(t)\,dt=\frac{1}{\Gamma(1-\alpha)}\int\limits_{0}^{T}\frac{d}{dt}\left(\int\limits_{t}^{T}(s-t)^{-\alpha}v(s)\,ds\right)dt.$
Using Fubini again, we obtain
$\frac{1}{\Gamma(1-\alpha)}\int\limits_{0}^{T}s^{-\alpha}v(s)\,ds=-I^{1-\alpha}_{0|T}v(0),$
and thus
$u(0)\int\limits_{0}^{T}\mathscr{D}_{t|T}^{\alpha}v(t)\,dt=-u(0)I^{1-\alpha}_{0|T}v(0).$
Substituting back, we find
$\int\limits_{0}^{T}v(t)\,\partial_{t}^{\alpha}u(t)\,dt=\int\limits_{0}^{T}(u(t)-u(0))\mathscr{D}_{t|T}^{\alpha}v(t)\,dt,$
which proves the result. ∎
2.1. Fundamental solution and Duhamel’s formula
A sufficiently smooth solution $u$ to (1.1) can be represented via Duhamel’s formula:
$u(t,x)$ $=\underbrace{\int_{\mathbb{R}^{N}}u_{0}(y)\,Z(t,x-y)\,dy}_{\mathcal{L}_{u_{0}}(t,x)}+\underbrace{\int_{0}^{t}\int_{\mathbb{R}^{N}}|u(\tau,y)|^{p}\,Y(t-\tau,x-y)\,dy\,d\tau}_{\mathcal{N}_{u}(t,x)}$ (2.7)
$\quad+\underbrace{\int_{0}^{t}\int_{\mathbb{R}^{N}}\tau^{\sigma}\,\mathbf{w}(y)\,Y(t-\tau,x-y)\,dy\,d\tau}_{\mathcal{S}_{\mathbf{w}}(t,x)}\;=\;\mathcal{G}u(t,x).$
Here $Z$ denotes the fundamental solution of the homogeneous equation
$\partial_{t}^{\alpha}u+(-\Delta)^{\mathsf{s}}u=0,$
while $Y=Z$ when $\alpha=1$, and $Y=\partial_{t}^{\alpha}Z$ for $\alpha\in(0,1)$; see [11, 24].
More precisely, both $Z$ and $Y$ admit self-similar representations of the form
$Z(t,x)=t^{-\frac{\alpha N}{2\mathsf{s}}}\,F\!\left(x\,t^{-\frac{\alpha}{2\mathsf{s}}}\right),\qquad Y(t,x)=t^{-\frac{\alpha N}{2\mathsf{s}}}\,G\!\left(x\,t^{-\frac{\alpha}{2\mathsf{s}}}\right),$
where $F$ and $G$ are the associated spatial profiles.
We next recall several fundamental properties of the kernels $Z$ and $Y$. The following key lemma gathers the main estimates and structural features that will be crucial in our analysis. For proofs and additional details, we refer to [15, 24, 26, 28, 11].
Lemma 2.2.
Let $Z$ and $Y$ be the spatial kernels associated with the linear part of equation (1.1), as defined in the integral formulation (2.7). The following properties hold
  1. (i)
    Both kernels exhibit the scaling forms
    $Z(t,x)=t^{-\frac{N\alpha}{2\mathsf{s}}}F\left(xt^{-\frac{\alpha}{2\mathsf{s}}}\right),\quad Y(t,x)=t^{-\left(1-\alpha+\frac{N\alpha}{2\mathsf{s}}\right)}G\left(xt^{-\frac{\alpha}{2\mathsf{s}}}\right),$
    where $F,G$ are positive, radially decreasing functions, smooth away from the origin.
  2. (ii)
    $\|Z(t,\cdot)\|_{L^{1}(\mathbb{R}^{N})}=1$ for all $t>0$.
  3. (iii)
    Let $q_{c}:=\frac{N}{N-2\mathsf{s}}$ if $N>2\mathsf{s}$, and $\infty$ if $N\leq 2\mathsf{s}$. For any $q\in[1,q_{c})$, $F\in L^{q}(\mathbb{R}^{N})$ and
    $\|Z(t,\cdot)\|_{L^{q}(\mathbb{R}^{N})}=t^{-\frac{N\alpha}{2\mathsf{s}}(1-\frac{1}{q})}\|F\|_{L^{q}(\mathbb{R}^{N})},\quad t>0.$ (2.8)
  4. (iv)
    For $t\geq t_{0}>0$,
    $\|Z(t,\cdot)-Z(t_{0},\cdot)\|_{L^{q}(\mathbb{R}^{N})}\leq C(t-t_{0})t_{0}^{-1-\frac{N\alpha}{2\mathsf{s}}(1-\frac{1}{q})},$
    where the constant depending only on $\alpha,\mathsf{s},N$, and q.
  5. (v)
    For any $\lambda>0$ and $q\in[1,\infty]$,
    $\|Z(t,\cdot)\|_{L^{q}(|x|>\lambda t^{\frac{\alpha}{2\mathsf{s}}})}=t^{-\frac{N\alpha}{2\mathsf{s}}(1-\frac{1}{q})}\|F\|_{L^{q}(|x|>\lambda)}.$
  6. (vi)
    For every $q\in[1,q_{c})$, the function $G$ belongs to $L^{q}(\mathbb{R}^{N})$, and the following identity holds:
    $\|Y(t,\cdot)\|_{L^{q}(\mathbb{R}^{N})}=t^{-(1-\alpha+\frac{N\alpha}{2\mathsf{s}}\left(1-\frac{1}{q}\right))}\|G\|_{L^{q}},\quad t>0.$ (2.9)
  7. (vii)
    For $0\leq\tau<t_{0}\leq t$,
    $\|Y(t-\tau,\cdot)-Y(t_{0}-\tau,\cdot)\|_{L^{q}(\mathbb{R}^{N})}\leq C(t-t_{0})(t_{0}-\tau)^{\alpha-\frac{N\alpha}{2\mathsf{s}}(1-\frac{1}{q})-2}.$ (2.10)
Remark 2.1.
It is straightforward to verify that
$1-\alpha+\frac{N\alpha}{2\mathsf{s}}\!\left(1-\frac{1}{q}\right)<1\quad\Longleftrightarrow\quad q<q_{c},$
which implies that
$Y(t,\cdot)\in L^{1}_{\mathrm{loc}}(0,\infty;L^{q})\quad\text{if and only if}\quad 1\leq q<q_{c}.$
A straightforward consequence of the kernel estimates (2.9) and (2.10) is the following estimate.
Lemma 2.3.
Let $\alpha\in(0,1)$ and $\gamma\in(0,1)$. There exists a constant $C>0$ such that for all $0<t_{0}\leq t$, the following Hölder-type estimate holds:
$\|Y(t,\cdot)-Y(t_{0},\cdot)\|_{L^{1}(\mathbb{R}^{N})}\leq C(t-t_{0})^{\gamma}t_{0}^{\alpha-1-\gamma}.$
Proof.
Let $0<t_{0}\leq t$ and $\gamma\in(0,1)$. From the time continuity estimate (2.10) with $\tau=0$ and $q=1$, we have
$\|Y(t,\cdot)-Y(t_{0},\cdot)\|_{L^{1}(\mathbb{R}^{N})}\leq C(t-t_{0})t_{0}^{\alpha-2}.$
Moreover, by the bound (2.9) with $q=1$, we get
$\|Y(t,\cdot)-Y(t_{0},\cdot)\|_{L^{1}(\mathbb{R}^{N})}\leq\|Y(t,\cdot)\|_{L^{1}}+\|Y(t_{0},\cdot)\|_{L^{1}}\leq C(t^{\alpha-1}+t_{0}^{\alpha-1})\leq Ct_{0}^{\alpha-1},$
where we used $t\geq t_{0}>0$ and $\alpha-1<0$. Interpolating between these two estimates yields
$\|Y(t,\cdot)-Y(t_{0},\cdot)\|_{L^{1}(\mathbb{R}^{N})}\leq C\left[(t-t_{0})t_{0}^{\alpha-2}\right]^{\gamma}\left[t_{0}^{\alpha-1}\right]^{1-\gamma}=C(t-t_{0})^{\gamma}t_{0}^{\alpha-1-\gamma}.$
For a parameter $\mathsf{s}\in(0,1)$ and a function $u:\mathbb{R}^{N}\to\mathbb{R}$, an equivalent definition of the fractional Laplacian of order $\mathsf{s}$ at a point $x\in\mathbb{R}^{N}$ is given by
$(-\Delta)^{\mathsf{s}}u(x)=C_{N,\mathsf{s}}\,\lim_{\varepsilon\to 0^{+}}\int_{\mathbb{R}^{N}\setminus B_{\varepsilon}(x)}\frac{u(x)-u(y)}{|x-y|^{N+2\mathsf{s}}}\,dy,$
whenever the limit exists. Here $B_{\varepsilon}(x)$ denotes the ball of radius $\varepsilon$ centered at $x$, and
$C_{N,\mathsf{s}}=\frac{2^{\,2\mathsf{s}-1}\,\Gamma\!\left(\frac{N}{2}+\mathsf{s}\right)}{\pi^{N/2}\,\Gamma(1-\mathsf{s})}$
is the normalization constant ensuring that $(-\Delta)^{\mathsf{s}}$ converges to the classical Laplacian as $\mathsf{s}\to 1^{-}$. Further equivalent formulations and additional properties of the fractional Laplacian can be found in [1, 30, 37].
We now recall a fundamental symmetry property that plays a central role in fractional integration by parts.
Lemma 2.4.
Let $\mathsf{s}\in(0,1)$. For any functions $u,v\in C^{2}(\mathbb{R}^{N})\cap L^{\infty}(\mathbb{R}^{N})$, the fractional Laplacian satisfies the identity
$\int_{\mathbb{R}^{N}}u(x)\,(-\Delta)^{\mathsf{s}}v(x)\,dx=\int_{\mathbb{R}^{N}}v(x)\,(-\Delta)^{\mathsf{s}}u(x)\,dx.$
We also recall the following form of the singular Gronwall inequality.
Lemma 2.5.
Let $T>0$, $\gamma>-1$, and $K>0$. Assume that $\varphi\in C([0,T];\mathbb{R}_{+})$ satisfies
$\varphi(t)\leq K\int_{0}^{t}(t-\tau)^{\gamma}\varphi(\tau)\,d\tau,\qquad\text{for all }t\in[0,T].$
Then $\varphi(t)=0$ for all $t\in[0,T]$.
We now establish global existence in Lebesgue spaces $L^{q}$ under suitable assumptions on the parameters $p,q$, together with a smallness condition.
Lemma 2.6.
Assume $\alpha,\mathsf{s}\in(0,1)$, $\alpha<\sigma<0$, and $2\mathsf{s}<N$. Suppose
$\frac{N\alpha-2\mathsf{s}\sigma}{N\alpha-2\mathsf{s}(\alpha+\sigma)}\leq p<\frac{N\alpha-2\mathsf{s}q\,\sigma}{N\alpha-2\mathsf{s}q\,(\alpha+\sigma)},$ (2.11)
and let $q$ belong to the admissible range
$p_{c}<q\leq p\,p_{c},\qquad q\geq p,$
where $p_{c}$ is given in (1.3). Let $u$ be a maximal $L^{q}$-solution of (1.1) corresponding to initial data $u_{0}\in L^{q}(\mathbb{R}^{N})$, and assume $\mathbf{w}\in L^{q}(\mathbb{R}^{N})$.
Then there exists a constant $\mathscr{M}>0$, depending only on $p,q,\alpha,\mathsf{s},N$, and $\sigma$, such that if
$t^{\,\frac{\alpha}{p-1}-\frac{N\alpha}{2\mathsf{s}q}}\Bigl(\|\mathcal{L}_{u_{0}}(t,\cdot)\|_{L^{q}}+\|\mathcal{S}_{\mathbf{w}}\|_{L^{q}}\Bigr)\leq\mathscr{M},\qquad\forall\,t>0,$ (2.12)
then the following a apriori bound holds:
$\sup_{\tau\in(0,t)}\tau^{\,\frac{\alpha}{p-1}-\frac{N\alpha}{2\mathsf{s}q}}\|u(\tau,\cdot)\|_{L^{q}}\leq 2\mathscr{M},\qquad\forall\,t\in(0,T_{\max}),$ (2.13)
where $T_{\max}\in(0,\infty]$ denotes the maximal existence time. In particular, $u$ is a global $L^{q}$-solution of (1.1).
Proof of Lemma 2.6.
First, we establish the existence of a maximum time $T_{\text{max}}>0$ and a maximal solution $u\in C\big([0,T_{\text{max}});L^{q}(\mathbb{R}^{N})\big)$ satisfying the integral equation (2.7). From the results of [11], we already know that $\mathcal{L}_{u_{0}},\mathcal{N}_{u}\in C([0,T];L^{q})$. To conclude that the full solution $u$ belongs to $C([0,T];L^{q})$, it remains to prove the continuity of the term $S_{\mathbf{w}}(t)$. We analyze the continuity at $t=0$ and for $t>0$ separately.
Let $\alpha\in(0,1)$, $\sigma+\alpha>0$, and $\mathbf{w}\in L^{q}(\mathbb{R}^{N})$. Using (2.9) and Young’s inequality, we get
$\|\mathcal{S}_{\mathbf{w}}(t)\|_{L^{q}}\leq\|\mathbf{w}\|_{L^{q}}\int\limits_{0}^{t}\tau^{\sigma}\|Y(t-\tau)\|_{L^{1}}ds\leq C\|\mathbf{w}\|_{L^{q}}\int\limits_{0}^{t}\tau^{\sigma}(t-\tau)^{\alpha-1}d\tau.$ (2.14)
The integral is equal to $t^{\sigma+\alpha}\mathscr{B}(\sigma+1,\alpha)$. Since $\sigma+\alpha>0$, we have
$\|\mathcal{S}_{\mathbf{w}}(t)\|_{L^{q}}\leq C\|\mathbf{w}\|_{L^{q}}t^{\sigma+\alpha}\to 0\quad\text{as }t\to 0^{+}.$
On the other hand, let $0\leq t_{1}<t_{2}\leq T$. We decompose
$\mathcal{S}_{\mathbf{w}}(t_{2})-\mathcal{S}_{\mathbf{w}}(t_{1})=\int\limits_{t_{1}}^{t_{2}}\tau^{\sigma}(Y(t_{2}-\tau)*\mathbf{w})d\tau+\int\limits_{0}^{t_{1}}\tau^{\sigma}\left(Y(t_{2}-\tau)-Y(t_{1}-\tau)\right)*\mathbf{w}d\tau.$
For the first term, an application of Minkowski’s inequality together with the kernel estimate from Lemma 2.2, while taking into account that $\tau\in[t_{1},t_{2}]\subset[0,T]$, yields
$\left\|\int\limits_{t_{1}}^{t_{2}}\tau^{\sigma}Y(t_{2}-\tau)*\mathbf{w}d\tau\right\|_{L^{q}}$ $\leq\|\mathbf{w}\|_{L^{q}}\int\limits_{t_{1}}^{t_{2}}\tau^{\sigma}\|Y(t_{2}-\tau)\|_{L^{1}}d\tau$ (2.15)
$\leq C\|\mathbf{w}\|_{L^{q}}\int\limits_{t_{1}}^{t_{2}}\tau^{\sigma}(t_{2}-\tau)^{\alpha-1}d\tau$ $\leq\frac{Ct_{1}^{\sigma}}{\alpha}\|\mathbf{w}\|_{L^{q}}(t_{2}-t_{1})^{\alpha}.$ (2.16)
We now turn to the second term. By applying Lemma 2.3 with the choices $_{0}=t_{1}-\tau, t=t$$t_{2}-\tau$, and with $\gamma$ restricted to $(0,\alpha+\sigma)$, we obtain
$\left\|\int\limits_{0}^{t_{1}}\tau^{\sigma}\left[Y(t_{2}-\tau,.)-Y(t_{1}-\tau,.)\right]*\mathbf{w}d\tau\right\|_{L^{q}}$ $\leq C_{\gamma}\|\mathbf{w}\|_{L^{q}}(t_{2}-t_{1})^{\gamma}\int\limits_{0}^{t_{1}}\tau^{\sigma}(t_{1}-\tau)^{\alpha-1-\gamma}d\tau$ (2.17)
$\leq C_{\gamma}T^{\alpha+\sigma-\gamma}\mathscr{B}(\sigma+1,\alpha-\gamma)\|\mathbf{w}\|_{L^{q}}(t_{2}-t_{1})^{\gamma}$
Combining (2.15) and (2.17), we conclude that
$\|\mathcal{S}_{\mathbf{w}}(t_{2})-\mathcal{S}_{\mathbf{w}}(t_{1})\|_{L^{q}}\leq C\|\mathbf{w}\|_{L^{q}}\left(t_{1}^{\sigma}|t_{2}-t_{1}|^{\alpha}+|t_{2}-t_{1}|^{\gamma}\right).$
where C depends on $\gamma$, $\alpha$, and $T$. This bound tends to zero as $|t_{2}-t_{1}|\to 0$, which establishes that $\mathcal{S}_{\mathbf{w}}\in C([0,T];L^{q})$. Since we have shown that $u\in C([0,T_{\max});L^{q})$, it follows that the limit $u(\cdot,T_{\max}):=\lim_{t\nearrow T_{\max}}u(\cdot,t)$ exists in the $L^{q}(\mathbb{R}^{N})$ .
Subsequently, we prove that if $T_{\text{max}}<\infty$, then $\limsup_{t\nearrow T_{\text{max}}}\|u(\cdot,t)\|_{L^{q}}=+\infty.$ We proceed by contradiction, assuming that $T:=T_{\text{max}}<\infty$ but that $\limsup_{t\nearrow T}\|u(\cdot,t)\|_{L^{q}}<+\infty,$ then we need to prove first an extension result that will characterize the behavior of the maximal solution at the boundary of it’s existence interval.
Let $h$ be such that $0<h<T_{\text{max}}/2$. For any $t\in(T_{max}-h,T_{max}+h)$, we define
$\mathcal{H}v(t,x)=$ $\underbrace{\int\limits_{\mathbb{R}^{N}}u_{0}(y)Z(t,x-y)dy}_{\mathcal{L}_{u_{0}}(t,x)}+\underbrace{\int\limits_{0}^{T_{max}-h}\int\limits_{\mathbb{R}^{N}}|u(\tau,y)|^{p}Y(t-\tau,x-y)dyd\tau}_{\mathcal{N}_{u}(t,x)}$
$+\underbrace{\int\limits_{T_{max}-h}^{t}\int\limits_{\mathbb{R}^{N}}|v(\tau,y)|^{p}Y(t-\tau,x-y)dyd\tau}_{\mathcal{F}_{v}(t,x)}+\underbrace{\int\limits_{0}^{t}\int\limits_{\mathbb{R}^{N}}\tau^{\sigma}\mathbf{w}(y)Y(t-\tau,x-y)dyd\tau}_{\mathcal{S}_{\mathbf{w}}(t,x)}.$
We establish a refined decomposition of $\mathcal{S}_{\mathbf{w}}(t)$ that depends on the position of the parameter $t$
$\mathcal{S}_{\mathbf{w}}(t)=\mathcal{S}_{1,\mathbf{w}}+\mathcal{S}_{2,\mathbf{w}}(t)+\mathcal{S}_{3,\mathbf{w}}(t)+\mathcal{S}_{4,\mathbf{w}}(t),$ (2.18)
where
$\mathcal{S}_{1,\mathbf{\mathbf{w}}}$ $=\int\limits_{0}^{T_{max}-h}\tau^{\sigma}\left(Y(T_{max}-\tau,.)*\mathbf{w}\right)d\tau,$
$\mathcal{S}_{2,\mathbf{\mathbf{w}}}(t)$ $=\int\limits_{0}^{T_{max}-h}\tau^{\sigma}\left(Y(t-\tau,.)-Y(T_{max}-\tau,.)\right)*\mathbf{w}d\tau,$
$\mathcal{S}_{3,\mathbf{\mathbf{w}}}(t)$ $=\int\limits_{T_{max}-h}^{\min(t,T_{max})}\tau^{\sigma}\left(Y(t-\tau,.)*\mathbf{w}\right)d\tau,$
$\mathcal{S}_{4,\mathbf{\mathbf{w}}}(t)$ $=\int\limits_{T_{max}}^{t}\tau^{\sigma}\left(Y(t-\tau,.)*\mathbf{w}\right)d\tau,\text{for}\quad t>T_{max}.$
We construct a solution $v$ that coincides with (2.1) on the interval $0,T_{\max}-h)$, where $\tau\in(0,T_{\max}/2)$ is sufficiently small. For positive parameters $h>0$ and $\delta>0$, we work in the space
$\mathcal{X}_{h,\delta}=\left\{v\in L^{\infty}\big((T_{max}-h,T_{max}+h);L^{q}(\mathbb{R}^{N})\big):\sup_{t\in[T_{max}-h,T_{max}+h]}\|v(t,\cdot)\|_{L^{q}}\leq\delta\right\}.$
Under the assumptions on $\sigma$ and $\alpha$, applying the kernel estimates (2.9), (2.10), for $Y$ and $Z$ together with Young’s inequality and the $\gamma$- Hölder regularity estimate (2.3) yields
$\|\mathcal{S}_{1,\mathbf{w}}\|_{L^{q}}$ $\leq\|\mathbf{w}\|_{L^{q}}\int\limits_{0}^{T_{max}-h}\tau^{\sigma}\|Y(T_{max}-\tau,.)\|_{L^{1}}d\tau$ (2.19)
$\leq C\|\mathbf{w}\|_{L^{q}}\int\limits_{0}^{T_{max}-h}\tau^{\sigma}(T_{max}-\tau)^{\alpha-1}d\tau\leq C\|\mathbf{w}\|_{L^{q}}T_{max}^{\sigma+\alpha}\mathscr{B}(\sigma+1,\alpha)=:\mathscr{K}_{1}.$
The estimate for $\mathcal{S}_{2,\mathbf{w}}$ follows using analogous techniques. Under the hypothesis $0<\gamma<\alpha$, Lemma 2.3 gives
$\|\mathcal{S}_{2,\mathbf{w}}\|_{L^{q}}$ $\leq\|\mathbf{w}\|_{L^{q}}\int\limits_{0}^{T_{max}-h}\tau^{\sigma}\|Y(t-\tau,.)-Y(.,T_{max}-\tau)\|_{L^{1}}d\tau$ (2.20)
$\leq C_{\gamma}\|\mathbf{w}\|_{L^{q}}h^{\gamma}\int\limits_{0}^{T_{max}-h}\tau^{\sigma}(T_{max}-\tau)^{\alpha-1-\gamma}d\tau$ (2.21)
$\leq C_{\gamma}\|\mathbf{w}\|_{L^{q}}h^{\gamma}T_{max}^{\sigma+\alpha-\gamma}\mathscr{B}(\sigma+1,\alpha-\gamma)=:\mathscr{K}_{2}h^{\gamma}.$ (2.22)
For $\tau\in[T_{max}-h,\min(t,T_{max})]$, we have $\tau\geq T_{max}-h\geq T_{max}/2$, so $\tau^{\sigma}\leq(T_{max}/2)^{\sigma}$
$\|\mathcal{S}_{3,\mathbf{w}}\|_{L^{q}}$ $\leq\|\mathbf{w}\|_{L^{q}}\int\limits_{T_{max}-h}^{\min(t,T_{max})}\tau^{\sigma}\|Y(t-\tau,.)\|_{L^{1}}d\tau$ (2.23)
$\leq C\|\mathbf{w}\|_{L^{q}}\int\limits_{T_{max}-h}^{\min(t,T_{max})}\tau^{\sigma}(t-\tau)^{\alpha-1}d\tau$ (2.24)
$C\|\mathbf{w}\|_{L^{q}}T_{max}^{\sigma}h^{\alpha}=:\mathscr{K}_{3}h^{\alpha}.$ (2.25)
Similarly, for $t>T_{max}$ and $\tau\in[T_{max},t]$, we have $\tau\geq T_{max}\geq T_{max}/2\Rightarrow\tau^{\sigma}\leq(T_{max}/2)^{\sigma}$, and
$\|\mathcal{S}_{4,\mathbf{w}}\|_{L^{q}}$ $\leq\|\mathbf{w}\|_{L^{q}}\int\limits_{T_{max}}^{t}\tau^{\sigma}\|Y(t-\tau,.)\|_{L^{1}}d\tau$ (2.26)
$\leq C\|\mathbf{w}\|_{L^{q}}\int\limits_{T_{max}}^{t}\tau^{\sigma}(t-\tau)^{\alpha-1}d\tau\leq CT_{max}^{\sigma}h^{\alpha}\|\mathbf{w}\|_{L^{q}}=:\mathscr{K}_{4}h^{\alpha}.$
We combine the estimates (2.19), (2.20), (2.24), and (2.26). By choosing $h$ small enough so that
$\mathscr{K}_{2}h^{\gamma}+(\mathscr{K}_{3}+\mathscr{K}_{4})h^{\alpha}\leq\frac{\mathscr{K}_{1}}{2}.$
Based on the estimates for $\mathcal{L}_{u_{0}}$, $\mathcal{N}_{u}$, and $\mathcal{F}_{v}$ obtained in [11], there exists, for all sufficiently small $\mathscr{K}$, a constant such that
$\|\mathcal{H}v\|_{L^{q}}\leq\|\mathcal{L}_{u_{0}}(t)+\mathcal{N}_{u}(t)+\mathcal{F}_{v}(t)\|_{L^{q}}+\|\mathcal{S}_{\mathbf{w}}\|_{L^{q}}\leq\mathscr{K}.$
Next, we establish the contractivity of $\mathcal{G}$ on $X_{h,\delta}$ for sufficiently small $h$. Specifically, for any $u,v\in X_{h,\delta}$ and $t\in[T_{\max}-h,T_{\max}+h]$,
$\|\mathcal{G}(u)(t)-\mathcal{G}(v)(t)\|_{L^{q}}$ $\leq\int\limits_{T_{\max}-h}^{t}\left\||u(\tau)|^{p-1}u(\tau)-|v(\tau)|^{p-1}v(\tau)\right\|_{L^{q/p}}\|Y(t-\tau,\cdot)\|_{L^{r}}\,d\tau$
$\leq C\int\limits_{T_{\max}-h}^{t}\|u(\tau)-v(\tau)\|_{L^{q}}\left(\|u(\tau)\|_{L^{q}}^{p-1}+\|v(\tau)\|_{L^{q}}^{p-1}\right)(t-\tau)^{\alpha-1-\frac{\alpha N}{2\mathsf{s}}\left(1-\frac{1}{r}\right)}\,d\tau$
$\leq 2C\delta^{p-1}\|u-v\|_{L^{\infty}(T_{\max}-h,T_{\max}+h;L^{q})}\int\limits_{T_{\max}-h}^{t}(t-\tau)^{\alpha-1-\frac{\alpha N}{2\mathsf{s}}\left(1-\frac{1}{r}\right)}\,d\tau$
$\leq C\delta^{p-1}h^{\alpha-\frac{\alpha N}{2\mathsf{s}}\left(\frac{p-1}{q}\right)}\|u-v\|_{L^{\infty}(T_{\max}-h,T_{\max}+h;L^{q})},$
Choosing $h$ small enough such that
$C\delta^{p-1}h^{\alpha-\frac{\alpha N}{2\mathsf{s}}\left(\frac{p-1}{q}\right)}\leq\frac{1}{2}.$
Hence, $\mathcal{G}$ is a contraction on $X_{h,\delta}$, and therefore admits a unique fixed point $v\in X_{h,\delta}$. Moreover, since $v(T_{\max})=\mathcal{G}(v)(T_{\max})=u(T_{\max}),$ the function $v$ defines the desired extension:
$v(t,x)=\begin{cases}u(t,x),&t\in[0,T_{\max}-h),\\ \tilde{u}(t,x),&t\in[T_{\max}-h,T_{\max}+h).\end{cases}$
Since at this stage $\|u\|_{L^{q}}$ remains bounded, we can extend the solution to the interval $0,T_{\max}+h)$. Assume, by contradiction, that $T_{\max}<\infty$ and
$\sup_{t\in[0,T_{\max})}\|u(t)\|_{L^{q}}\leq M<\infty.$
From the continuity result established earlier, the mapping $t\mapsto u(\cdot,t)$ is continuous in $L^{q}(\mathbb{R}^{N})$ on $0,T_{\max})$. Therefore, the limit
$u(T_{\max},\cdot):=\lim_{t\to T_{\max}^{-}}u(t,\cdot)$
exists in $L^{q}(\mathbb{R}^{N})$, and in particular, $u(\cdot,T_{\max})\in L^{q}(\mathbb{R}^{N})$.
By the local existence and extension result for mild solutions, if the solution $u$ belongs to $L^{q}(\mathbb{R}^{N})$ at time $T_{\max}$, then it can be extended to a larger time interval $0,T_{\max}+h)$, for some $h>0$, while remaining a solution in the same class. This contradicts the maximality of $T_{\max}$. Hence, the initial assumption is false, and we must have:
$\text{If }T_{\max}<\infty,\text{ then }\limsup_{t\to T_{\max}^{-}}\|u(t)\|_{L^{q}}=+\infty.$ (2.27)
On the other hand, we have
$\sup_{\tau\in(0,t)}\tau^{\beta}\|u(\tau,\cdot)\|_{L^{q}}$ $\leq\sup_{\tau\in(0,t)}\tau^{\beta}\|\mathcal{L}_{u_{0}}\|_{L^{q}}+\sup_{\tau\in(0,t)}\tau^{\beta}\|\mathcal{N}_{u}\|_{L^{q}}+\mathscr{M}_{2}\sup_{\tau\in(0,t)}\tau^{\beta+\sigma+\alpha}$
$\leq\mathscr{M}_{0}+\mathscr{M}_{1}\left(\sup_{\tau\in(0,t)}\|\tau^{\beta}u(\tau,\cdot)\|_{L^{q}}\right)^{p}+\mathscr{M}_{2}t^{\beta+\sigma+\alpha}$
where $\mathscr{M}_{1}=\|G\|_{L^{r}}\int\limits_{0}^{1}\tau^{-\beta p}(1-\tau)^{-(1-\alpha+\frac{N\alpha}{2\mathsf{s}q}(p-1))}d\tau$, and $\mathscr{M}_{2}=\|\mathbf{w}\|_{L^{q}}\|G\|_{L^{1}}\mathscr{B}(\sigma+1,\alpha).$
Under the conditions (2.11) and $0<\beta<\frac{1}{p}$, which ensure the convergence of the integral $\mathscr{M}_{1}$, and under the hypothesis $\sigma>-\alpha$, we obtain a result that, through a bootstrap argument, yields
$\sup_{\tau\in(0,t)}\|\tau^{\beta}u(\tau,\cdot)\|_{L^{q}}\leq 2\mathscr{M}\quad\text{for all}\quad t\in(0,T).$
Therefore, we deduce that
$\|u(t,\cdot)\|_{L^{q}(\mathbb{R}^{N})}\leq 2\mathscr{M}t^{-\beta}\quad\text{for all }t\in(0,T).$
In particular, for any $\tau\in(0,T)$
$\sup_{t\in[0,\tau]}\|u(t,\cdot)\|_{L^{q}(\mathbb{R}^{N})}\leq\max\left\{\|u_{0}\|_{L^{q}(\mathbb{R}^{N})},\ \sup_{\tau\in(0,T)}2\mathscr{M}\tau^{-\beta}\right\}<\infty,$
which contradicts (2.27). Therefore, $T_{\max}=\infty$, completing the proof. ∎
3. Proofs of the main results
3.1. Proof of Theorem 1.1
For given $T>0$, and $u_{0},\mathbf{w}\in L^{q}(\mathbb{R}^{N})$, and let $\delta$ be a sufficiently small positive constant. We define the function space
$Y=\left\{u\in C\big((0,T),L^{q}(\mathbb{R}^{N})\big),\left\|u\right\|_{L^{\infty}((0,T),L^{q}({\mathbb{R}^{N}})}\leq\delta\right\},$
equipped with the metric $d(u,v)=\sup_{t\in[0,T]}\|u-v\|_{L^{q}(\mathbb{R}^{N})}$, for $u,v\in Y$. Since $(Y,d)$ is a complete metric space.
Let $\alpha,\mathsf{s}\in(0,1),$ $p>1$, $\sigma>-1$, suppose $u_{0},\mathbf{w}\in L^{q}(\mathbb{R}^{N})$ satisfy $\|u_{0}\|_{L^{q}}+\|\mathbf{w}\|_{L^{q}}<\varepsilon$ for some $\varepsilon>0$. We define the operator $\mathcal{G}$ on $Y$ by
$\begin{split}\mathcal{G}u(t,x)&=\int\limits_{\mathbb{R}^{N}}u_{0}(y)Z(t,x-y)dy+\int\limits_{0}^{t}\int\limits_{\mathbb{R}^{N}}|u(\tau,y)|^{p}Y(t-\tau,x-y)dyd\tau\\ &+\int\limits_{0}^{t}\int\limits_{\mathbb{R}^{N}}\tau^{\sigma}\mathbf{w}(y)Y(t-\tau,x-y)dyd\tau.\end{split}$ (3.1)
Suppose that $u\in C([0,T],L^{q}(\mathbb{R}^{N}))$ with $q>p_{c}$. It is then straightforward to verify that (3.1) also belongs to $C([0,T],L^{q}(\mathbb{R}^{N}))$.
Let $u_{0},\mathbf{w}\in L^{q}(\mathbb{R}^{N})$. Using the estimates for the kernels $Z$ and $Y$ from Lemma 2.2, for $t\in[0,T]$, we obtain
$\|\mathcal{G}u(t)\|_{L^{q}}$ $\leq\|u_{0}\|_{L^{q}(\mathbb{R}^{N})}+\left\|\int\limits_{0}^{t}\int\limits_{\mathbb{R}^{N}}Y(t-\tau,x-y)|u(\tau,y)|^{p}dyd\tau\right\|_{L^{q}(\mathbb{R}^{N})}$
$+\left\|\int\limits_{0}^{t}\int\limits_{\mathbb{R}^{N}}\tau^{\sigma}Y(t-\tau,x-y)\mathbf{w}(y)dyd\tau\right\|_{L^{q}(\mathbb{R}^{N})}$
$\leq\|u_{0}\|_{L^{q}}+C\delta^{p}\int\limits_{0}^{t}(t-\tau)^{-\frac{\alpha N}{2\mathsf{s}}\left(\frac{p-1}{q}\right)+\alpha-1}d\tau+C\|\mathbf{w}\|_{L^{q}(\mathbb{R}^{N})}\int\limits_{0}^{t}\tau^{\sigma}(t-\tau)^{\alpha-1}d\tau.$
When $N\geq 2\mathsf{s}$, the condition $q>p_{c}$ ensures that the exponent satisfies $\alpha-\frac{\alpha N}{2\mathsf{s}}\left(\frac{p-1}{q}\right)-1>-1$, which guarantees the convergence of the time integral. For $N<2\mathsf{s}$, this condition is automatically satisfied for $q\geq p$. For $t\in(0,T)$, the change of variable $\tau=t\theta$ gives
$\|\mathcal{G}u(t)\|_{L^{q}(\mathbb{R}^{N})}$ $\leq\|u_{0}\|_{L^{q}(\mathbb{R}^{N})}+C\delta^{p}t^{\alpha-\frac{\alpha N}{2\mathsf{s}}\left(\frac{p-1}{q}\right)}\int\limits_{0}^{1}(1-\theta)^{\alpha-\frac{\alpha N}{2\mathsf{s}}\left(\frac{p-1}{q}\right)-1}d\theta$
$+Ct^{\alpha+\sigma}\|\mathbf{w}\|_{L^{q}(\mathbb{R}^{N})}\int\limits_{0}^{1}\theta^{\sigma}(1-\theta)^{\alpha-1}d\theta$
$\leq\|u_{0}\|_{L^{q}(\mathbb{R}^{N})}+C\delta^{p}T^{\alpha-\frac{\alpha N}{2\mathsf{s}}\left(\frac{p-1}{q}\right)}+CT^{\alpha+\sigma}\mathscr{B}\left(\sigma+1,\alpha\right)\|\mathbf{w}\|_{L^{q}(\mathbb{R}^{N})}.$
Hence, we can choose $T$ small enough such that
$C\delta^{p}T^{\alpha-\frac{\alpha N}{2\mathsf{s}}\left(\frac{p-1}{q}\right)}+CT^{\alpha+\sigma}\mathscr{B}(\sigma+1,\alpha)\|\mathbf{w}\|_{L^{q}(\mathbb{R}^{N})}\leq\delta-\|u_{0}\|_{L^{q}}$
This shows that $\mathcal{G}$ is a mapping from $Y$ into itself. It remains to demonstrate that the self-mapping $\mathcal{G}:Y\rightarrow Y$ is a contraction to complete the proof. Applying similar arguments as mentioned above, we can establish that for functions $u$ and $v$ in $Y$
$\|\mathcal{G}(u)-\mathcal{G}(v)\|_{L^{q}(\mathbb{R}^{N})}$ $\leq C\int\limits_{0}^{t}(t-\tau)^{\alpha-1-\frac{\alpha N}{2\mathsf{s}}\left(\frac{p-1}{q}\right)}\||u|^{p}-|v|^{p}\|_{L^{q/p}(\mathbb{R}^{N})}\,ds$
$\leq C\int\limits_{0}^{t}(t-\tau)^{\alpha-1-\frac{\alpha N}{2\mathsf{s}}\left(\frac{p-1}{q}\right)}\left(\|u\|^{p-1}_{L^{q}}+\|v\|^{p-1}_{L^{q}}\right)\|u-v\|_{L^{q}(\mathbb{R}^{N})}\,d\tau.$
where, for $p>1$, we have used the inequality
$\left||u|^{p}-|v|^{p}\right|\leq C|u-v|\left(|u|^{p-1}+|v|^{p-1}\right).$
It follows that,
$\|\mathcal{G}(u)-\mathcal{G}(v)\|_{L^{q}(\mathbb{R}^{N})}$ $\leq C\delta^{p-1}\int\limits_{0}^{t}(t-\tau)^{\alpha-1-\frac{\alpha N}{2\mathsf{s}}\left(\frac{p-1}{q}\right)}\,d\tau\|u-v\|_{Y}$
$=C\delta^{p-1}T^{\alpha-\frac{\alpha N}{2\mathsf{s}}\left(\frac{p-1}{q}\right)}\int\limits_{0}^{1}(1-\tau)^{\alpha-1-\frac{\alpha N}{2\mathsf{s}}\left(\frac{p-1}{r}\right)}\,d\tau\|u-v\|_{Y}.$
By the Banach fixed point theorem, there exists a unique $u\in Y$ such that $u=\mathcal{G}u$, which is the desired solution. The uniqueness is proven using a standard Gronwall argument. Consider two mild solutions $u,v\in C([0,T],L^{q}(\mathbb{R}^{N}))$ of (1.1). Under the parameter constraints $q>p_{c}$ and $q\geq p$, the following key estimate holds:
$\|\mathcal{G}u(t)-\mathcal{G}v(t)\|_{L^{q}(\mathbb{R}^{N})}$ $=\|u(t)-v(t)\|_{L^{q}(\mathbb{R}^{N})}$
$\leq CM\int\limits_{0}^{t}(t-\tau)^{\alpha-1-\frac{\alpha N(p-1)}{2sq}}\|u(\tau)-v(\tau)\|_{L^{q}(\mathbb{R}^{N})}d\tau,$
where $M=\sup_{t\in[0,T]}\left(\|u(t)\|_{L^{q}}^{p-1}+\|v(t)\|_{L^{q}}^{p-1}\right)$ is a finite constant. Since the exponent
$\alpha-1-\frac{\alpha N(p-1)}{2\mathsf{s}q}>-1,$
the conditions for the fractional Gronwall inequality (2.5) are satisfied , from which we conclude $u\equiv v$ on $[0,T]$.
3.2. Proof of Theorem 1.2
Let $u$ be a global weak solution to problem (1.1), so that $u\in(0,\infty)$. For given positive numbers $T$ and $R$ satisfying $TR^{2\mathsf{s}/\alpha}<\infty$, we can choose $T^{*}>TR^{2\mathsf{s}/\alpha}$.
To establish a contradiction, we proceed with a carefully chosen test function argument. We begin by introducing a smooth cutoff function $\psi\in\mathbf{C}^{2}(\mathbb{R}^{N})$ with the following properties
$0\leqslant\psi\leqslant 1\quad\text{and}\quad\psi(z)=\begin{cases}1&\text{if }|z|\leq 1,\\ 0&\text{if }|z|\geq 2.\end{cases}$
The core of our argument relies on the rescaled test function
$\psi_{R}(t,x)=\psi\left(\frac{t}{R^{2\mathsf{s}/\alpha}}\right)\psi\left(\frac{|x|}{R}\right),\quad R>0.$
By multiplying problem (1.1) by $\psi_{R}$, applying the integration by parts formula from Lemmas (2.5) and (2.4), we exploiting the properties of $\psi$, we derive the desired contradiction.
$\int\limits_{0}^{T}\int\limits_{\mathbb{R}^{N}}u_{0}(x)\,\mathscr{D}_{t|T}^{\alpha}\psi_{R}(t,x)\,dx\,dt+\int\limits_{0}^{T}\int\limits_{\mathbb{R}^{N}}|u|^{p}\psi_{R}(t,x)\,dx\,dt+\int\limits_{0}^{T}\int\limits_{\mathbb{R}^{N}}t^{\sigma}\mathbf{w}\,\psi_{R}(t,x)\,dx\,dt$ (3.2)
$=\int\limits_{0}^{T}\int\limits_{\mathbb{R}^{N}}u(-\Delta)^{s}\psi_{R}(t,x)\,dx\,dt+\int\limits_{0}^{T}\int\limits_{\mathbb{R}^{N}}u\,\mathscr{D}_{t|T}^{\alpha}\psi_{R}(t,x)\,dx\,dt$
For our analysis, we define the following key integral quantities $I=\int\limits_{0}^{T}\int\limits_{\mathbb{R}^{N}}|u||(-\Delta)^{\mathsf{s}}\psi_{R}(t,x)|\,dx\,dt,$ and $J=\int\limits_{0}^{T}\int\limits_{\mathbb{R}^{N}}|u||\mathscr{D}_{t|T}^{\alpha}\psi_{R}(t,x)|\,dx\,dt$. To derive suitable bounds for both $I$ and $J$, we apply the $\varepsilon$-Young inequality and obtain the estimates
$I$ $\leq\frac{1}{2}\int\limits_{0}^{T}\int\limits_{\mathbb{R}^{N}}|u|^{p}\psi_{R}(t,x)\,dx\,dt+C\int\limits_{0}^{TR^{\frac{2s}{\alpha}}}\psi\left(\frac{t}{R^{2\mathsf{s}/\alpha}}\right)\,dt\int\limits_{\mathbb{R}^{N}}\left|(-\Delta)^{\mathsf{s}}\psi\left(\frac{|x|}{R}\right)\right|^{\frac{p}{p-1}}\left|\psi\left(\frac{|x|}{R}\right)\right|^{-\frac{1}{p-1}}\,dx$
$\leq\frac{1}{2}\int\limits_{0}^{T}\int\limits_{\mathbb{R}^{N}}|u|^{p}\psi_{R}(t,x)\,dx\,dt+CR^{N+\frac{2\mathsf{s}}{\alpha}-\frac{2\mathsf{s}p}{p-1}}.$
Parallel reasoning applied to the second integral, using the Riemann-Liouville fractional derivative definition (2.2), produces
$J$ $\leq\frac{1}{2}\int\limits_{0}^{T}\int\limits_{\mathbb{R}^{N}}|u|^{p}\psi_{R}(t,x)\,dx\,dt+\int\limits_{0}^{TR^{\frac{2\mathsf{s}}{\alpha}}}\left|\mathscr{D}_{t|TR^{\frac{\mathsf{s}}{\alpha}}}^{\alpha}\psi\left(\frac{t}{R^{\frac{2\mathsf{s}}{\alpha}}}\right)\right|^{\frac{p}{p-1}}\psi\left(\frac{t}{R^{2\mathsf{s}/\alpha}}\right)^{-\frac{1}{p-1}}\,dt\int\limits_{\mathbb{R}^{N}}\psi\left(\frac{|x|}{R}\right)\,dx$
$\leq\frac{1}{2}\int\limits_{0}^{T}\int\limits_{\mathbb{R}^{N}}|u|^{p}\psi_{R}(t,x)\,dx\,dt+CR^{N+\frac{2\mathsf{s}}{\alpha}-\frac{2\mathsf{s}p}{p-1}}.$
Since $\psi$ is decreasing, the function $t\mapsto\psi\left(\frac{t}{R^{2\mathsf{s}/\alpha}}\right)$ is decreasing in $t$, hence $\mathscr{D}_{t|T}^{\alpha}\psi_{R}(t,x)\geq 0$ for $T$ sufficiently large. With $u_{0}(x)\geq 0$, the integrand is nonnegative, so
$\int\limits_{0}^{T}\int\limits_{\mathbb{R}^{N}}u_{0}(x)\,\mathscr{D}_{t|T}\psi_{R}(t,x)\,dx\,dt\geq 0,$
together with the previously established estimates, implies that
$\int\limits_{0}^{TR^{\frac{2\mathsf{s}}{\alpha}}}t^{\sigma}\psi\left(\frac{t}{R^{\frac{2\mathsf{s}}{\alpha}}}\right)\,dt\int\limits_{\mathbb{R}^{N}}\mathbf{w}(x)\psi\left(\frac{|x|}{R}\right)\,dx\leqslant CR^{N+\frac{2\mathsf{s}}{\alpha}-\frac{2\mathsf{s}p}{p-1}}$
On the other hand, we take into account that $\mathbf{w}$ is in $L^{1}$. Then, for $R$ sufficiently large, we obtain by the Dominated Convergence Theorem that $\int\limits_{\mathbb{R}^{N}}\mathbf{w}(x)\psi\left(\frac{|x|}{R}\right)dx\longrightarrow\int\limits_{\mathbb{R}^{N}}\mathbf{w}(x)dx\geq 0$ Moreover, we can bound the second term on the right-hand side of (3.2) from below as follows
$\int\limits_{0}^{TR^{\frac{2\mathsf{s}}{\alpha}}}\int\limits_{\mathbb{R}^{N}}\mathbf{w}(x)\psi_{R}(t,x)dxdt$ $\geqslant\int\limits_{\frac{TR^{\frac{2\mathsf{s}}{\alpha}}}{2}}^{TR^{\frac{2\mathsf{s}}{\alpha}}}t^{\sigma}\psi\left(\frac{t}{R^{\frac{2\mathsf{s}}{\alpha}}}\right)dt\int\limits_{\mathbb{R}^{N}}\psi\left(\frac{|x|}{R}\right)\mathbf{w}(x)\,dx$
$\geq R^{\frac{2\mathsf{s}}{\alpha}(\sigma+1)}\int\limits_{\frac{1}{2}}^{1}\tau^{\sigma}\psi\left(\tau\right)d\tau\int\limits_{\mathbb{R}^{N}}\psi\left(\frac{|x|}{R}\right)\mathbf{w}(x)\,dx.$
We then have the estimate,
$\int_{\mathbb{R}^{N}}\mathbf{w}(x)\,dx\leq CR^{\gamma},\text{where}\quad\gamma=N-\frac{2\mathsf{s}p}{p-1}-\frac{2\mathsf{s}\sigma}{\alpha}.$
Observing that $1<p<\frac{N\alpha-2\mathsf{s}\sigma}{N\alpha-2\mathsf{s}(\alpha+\sigma)}$ and letting $R\to\infty$, we obtain $\int_{\mathbb{R}^{N}}\mathbf{w}(x)dx\leq 0,$ which is clearly a contradiction. This completes the proof.
3.3. Proof of Theorem 1.3
The global existence result of the theorem in the case without a forcing term has been studied in [21]. In that setting, the authors establish their result by employing a contradiction argument combined with suitable test functions and scaling techniques, through which they identify the critical exponents governing global existence.
In this section, we extend the analysis to include the presence of a forcing term. The critical exponent $p_{c}$ defined in (1.3), together with the exponent $r_{c}$ appearing in Theorem 1.3, plays a fundamental role in the development of the global existence analysis presented here.
We can select a positive constant $q$ that satisfies:
$\max\left(\frac{2\mathsf{s}}{Np(p-1)},\frac{2\mathsf{s}\alpha+2s\sigma(p-1)}{N\alpha(p-1)}\right)<\frac{1}{q}<\min\left(\frac{2\mathsf{s}}{N(p-1)},\frac{N\alpha-2\mathsf{s}(\alpha+\sigma)}{N\alpha-2\mathsf{s}\sigma}\right),\quad q\geq p$ (3.3)
Moreover, the assumptions $\alpha\in(0,1)$, $p>p^{*}=\frac{N\alpha-2\mathsf{s}\sigma}{N\alpha-2\mathsf{s}(\alpha+\sigma)}$ and $\alpha<\sigma<0$ imply that $p_{c},{r_{c}}>1$. We now establish a sufficient condition for the existence of a global solution. Specifically, we prove that there exists a constant $\epsilon>0$, depending solely on $p,q,\alpha,\sigma$, and $N$, such that for any initial data $u_{0}\in L^{p_{c}}$ and $\mathbf{w}\in L^{r_{c}}$ satisfying the smallness condition $\|u_{0}\|_{L^{p_{c}}}+\|\mathbf{w}\|_{L^{r_{c}}}<\epsilon$, a global solution exists. Here, $q$ is considered a fixed parameter, and we define
$\beta-\frac{N\alpha}{2\mathsf{s}}\left(\frac{1}{p_{c}}-\frac{1}{q}\right)=\alpha-\beta(p-1)-\frac{\alpha N}{2\mathsf{s}}\left(\frac{p-1}{q}\right)=\beta+\alpha-\frac{N\alpha}{2\mathsf{s}}\left(\frac{1}{{r_{c}}}-\frac{1}{q}\right)+\sigma=0$
$\beta=\frac{\alpha}{p-1}-\frac{\alpha N}{2q\mathsf{s}}$
It is straightforward to verify from condition (3.3) that $0<\beta p\leq\alpha<1$, where the last inequality is strict when $\alpha\in(0,1)$. Since $q\geq p$, this further implies $0<\beta\leq\frac{1}{q}$, which is the required bound. Now, let $\delta$ be a sufficiently small positive constant. We define the function space $Y$ as follows
$Y=\left\{u\in L_{Loc}^{\infty}((0,\infty),L^{q}(\mathbb{R}^{N}))\;;\>\sup_{t>0}t^{\beta}\|u(t)\|_{L^{q}}\leqslant\delta\right\}$
with the distance defined as $d(u,v)=\sup_{t>0}t^{\beta}\|u-v\|_{L^{q}}$, it is evident that $(Y,d)$ forms a complete metric space. We will demonstrate that the equation
$\mathcal{G}u(t,x)=$ $\int\limits_{\mathbb{R}^{N}}u_{0}(y)Z(t,x-y)dy+\int\limits_{0}^{t}\int\limits_{\mathbb{R}^{N}}|u(y,\tau)|^{p}Y(t-\tau,x-y)dyd\tau$
$+\int\limits_{0}^{t}\int\limits_{\mathbb{R}^{N}}\tau^{\sigma}\mathbf{w}(y)Y(t-\tau,x-y)dyd\tau.$
has a unique fixed point in $Y$. Assuming $u_{0}\in L^{p_{c}}$ and using Lemma 2.2, we select $r$ satisfying $1-\frac{1}{r}=\frac{1}{p_{c}}-\frac{1}{q}$. From condition (3.3), we observe that $1-\frac{1}{r}\in\big(0,\frac{2\mathsf{s}}{N}\big)$, which implies
$t^{\beta}\|\mathcal{L}_{u_{0}}(t,\cdot)\|_{L^{q}(\mathbb{R}^{N})}$ $\leq\|u_{0}\|_{L^{p_{c}}(\mathbb{R}^{N})}\|Z(t,\cdot)\|_{L^{r}(\mathbb{R}^{N})}$
$\leq t^{\beta-\frac{N\alpha}{2\mathsf{s}}\left(1-\frac{1}{r}\right)}\|u_{0}\|_{L^{p_{c}}}\|F\|_{L^{r}}:=\eta\|F\|_{L^{r}}.$
In the same manner, we choose $m$ such that $\frac{1}{m}=1-\frac{p}{q}+\frac{1}{q}.$ Since $p\geq 1$ and $q>\frac{N\alpha}{N\alpha-2\mathsf{s}(\alpha+\sigma)}$, we have $0\leq 1-\frac{1}{m}<\frac{2\mathsf{s}}{N}$, or equivalently, $m\in[1,\frac{N}{N-2s\mathsf{s}})$. Hence, (2.2) applies and yields the estimate $1-\alpha+\frac{N\alpha}{2\mathsf{s}}\left(1-\frac{1}{m}\right)<1.$ Thus, for $0\leq t\leq t_{0}$,
$\left\|\mathcal{N}(u)(t,\cdot)\right\|_{L^{q}(\mathbb{R}^{N})}$ $=\int\limits_{0}^{t}\left\||u(s,\cdot)|^{p}\right\|_{L^{\frac{q}{p}}(\mathbb{R}^{N})}\left\|Y(t-\tau,\cdot)\right\|_{L^{m}(\mathbb{R}^{N})}d\tau$
$\leq\left\|G\right\|_{L^{m}(\mathbb{R}^{N})}\sup_{\tau>0}\tau^{\beta p}\left\|u(\tau,\cdot)\right\|_{L^{q}(\mathbb{R}^{N})}^{p}\int\limits_{0}^{t}\tau^{-\beta p}(t-\tau)^{-(1-\alpha+\frac{N\alpha}{2\mathsf{s}q}(p-1))}d\tau$
$=\delta^{p}t^{\alpha-\frac{N\alpha}{2\mathsf{s}q}(p-1)-\beta p}\int\limits_{0}^{1}\tau^{-\beta p}(1-\tau)^{-(1-\alpha+\frac{N\alpha}{2\mathsf{s}q}(p-1))}d\tau\left\|G\right\|_{L^{m}(\mathbb{R}^{N})}$
$=\delta^{p}t^{-\beta}\mathscr{B}\left(1-\beta p,\alpha-\frac{\alpha N}{2\mathsf{s}}\left(\frac{p-1}{q}\right)\right)\left\|G\right\|_{L^{m}(\mathbb{R}^{N})}$
Here, $\mathscr{B}$ represents the beta function. We observe that by (3.3) and the fact that $\beta p\leqslant\beta q<1$, $\mathscr{B}$ is well-defined.
Proceeding similarly, we define $1+\frac{1}{q}=\frac{1}{{r_{c}}}+\frac{1}{\varrho}$. The condition on $q$ given in (3.3) ensures that $1-\frac{1}{\varrho}\in(0,\frac{2\mathsf{s}}{N})$. Given that $\sigma>-1$, we may use lemma (2.2) to obtain
$\|\mathcal{S}_{\mathbf{w}}(t,\cdot)\|_{L^{q}(\mathbb{R}^{N})}$ $\leq\|\mathbf{w}\|_{L^{{r_{c}}}(\mathbb{R}^{N})}\|G\|_{L^{\varrho}}\int\limits_{0}^{t}\tau^{\sigma}(t-\tau)^{-(1-\alpha+\frac{N\alpha}{2\mathsf{s}}(1-\frac{1}{\varrho}))}d\tau$
$\leq t^{\alpha-\frac{N\alpha}{2\mathsf{s}}(1-\frac{1}{\varrho})+\sigma}\|\mathbf{w}\|_{L^{{r_{c}}}(\mathbb{R}^{N})}\|G\|_{L^{\varrho}}\int\limits_{0}^{1}\tau^{\sigma}(1-\tau)^{\alpha-1-\frac{N\alpha}{2\mathsf{s}}(1-\frac{1}{\varrho})}d\tau$
$=t^{-\beta}\mathscr{B}\left(\sigma+1,\alpha-\frac{N\alpha}{2\mathsf{s}}(1-\frac{1}{\varrho})\right)\|\mathbf{w}\|_{L^{{r_{c}}}(\mathbb{R}^{N})}\|G\|_{L^{\varrho}(\mathbb{R}^{N})}.$
Therefore, based on the estimates above, choosing sufficiently small $\epsilon$ and $\delta>0$ leads to the conclusion that
$\sup_{t>0}t^{\beta}\|(\mathcal{G}u)(t)\|_{L^{r}}\leq C\left(\|u_{0}\|_{L^{p_{c}}}+\delta^{p}+\|\mathbf{w}\|_{L^{{r_{c}}}}\right)\leqslant C\epsilon$
This shows that $\mathcal{G}$ is a mapping from $Y$ into itself. It remains to demonstrate that the self-mapping $\mathcal{G}:Y\rightarrow Y$ is a contraction to complete the proof. By applying similar arguments as mentioned above, we can establish that for functions $u$ and $v$ in $L_{Loc}^{\infty}((0,\infty),L^{q}(\mathbb{R}^{N}))$
$t^{\beta}\|\mathcal{G}(u)-\mathcal{G}(v)\|_{L^{q}(\mathbb{R}^{N})}$ $\leq Ct^{\beta}\|G\|_{L^{r}}\int\limits_{0}^{t}(t-\tau)^{\alpha-1-\frac{\alpha N}{2\mathsf{s}}\left(\frac{p-1}{q}\right)}\|u^{p}-v^{p}\|_{L^{\frac{q}{p}}(\mathbb{R}^{N})}\,d\tau$
$\leq C\|G\|_{L^{r}}t^{\beta}\int\limits_{0}^{t}(t-\tau)^{\alpha-1-\frac{\alpha N}{2\mathsf{s}}\left(\frac{p-1}{q}\right)}\left(\|u\|^{p-1}_{L^{q}}+\|v\|^{p-1}_{L^{q}}\right)\|u-v\|_{L^{q}(\mathbb{R}^{N})}\,d\tau$
$\leq Ct^{\beta}\|G\|_{L^{r}}\delta^{p-1}\int\limits_{0}^{t}(t-\tau)^{\alpha-1-\frac{\alpha N}{2\mathsf{s}}\left(\frac{p-1}{q}\right)}\tau^{-p\beta}\,d\tau\|u-v\|_{Y}$
$=C\delta^{p-1}t^{\beta-p\beta-\alpha-\frac{\alpha N}{2\mathsf{s}}\left(\frac{p-1}{q}\right)}\int\limits_{0}^{1}(1-\tau)^{\alpha-1-\frac{\alpha N}{2\mathsf{s}}\left(\frac{p-1}{q}\right)}\tau^{-p\beta}\,d\tau\|u-v\|_{Y}$
$=C\delta^{p-1}\mathscr{B}\left(1-p\beta,\alpha-\frac{\alpha N}{2\mathsf{s}}\left(\frac{p-1}{q}\right)\right)\|u-v\|_{Y}.$
Hence, $\mathcal{G}$ is clearly a contraction on $Y$. Thus, by the contraction mapping principle, $\mathcal{G}$ has a unique fixed point $u$ in $Y$ which is the desired solution.
3.4. Proof of Theorem 1.4
The proof consists of verifying that the following inequality from Lemma 2.6 holds:
$t^{\beta}\left(\|\mathcal{L}_{u_{0}}(t,\cdot)\|_{L^{q}}+\|\mathcal{S}_{\mathbf{w}}(t,\cdot)\|_{L^{q}}\right)\leq\mathscr{M}\quad\text{for all }t>0,$
where $\beta=\frac{\alpha}{p-1}-\frac{N\alpha}{2\mathsf{s}q}.$ We proceed considering cases $t>R_{0}$ and $t\leq R_{0}$ separately. For $t\leq R_{0}$, we exploit the contraction properties of the kernels $Z$ and $Y$ given in Lemma 2.2 to obtain $\|\mathcal{L}_{u_{0}}(t,\cdot)\|_{L^{q}}\leq\|u_{0}\|_{L^{q}}$. For the source term, Young’s inequality and the bound $\|Y(t,\cdot)\|_{L^{1}}\leq Ct^{\alpha-1}$ yield
$\|\mathcal{S}_{\mathbf{w}}(t,\cdot)\|_{L^{q}}\leq\|\mathbf{w}\|_{L^{q}}\int\limits_{0}^{t}\tau^{\sigma}\|Y(t-\tau,\cdot)\|_{L^{1}}d\tau\leq C\|\mathbf{w}\|_{L^{q}}t^{\alpha+\sigma}.$
Choosing
$R_{0}=\min\left\{\left(\frac{\mathscr{M}}{2\|u_{0}\|_{L^{q}}}\right)^{1/\beta},\left(\frac{\mathscr{M}}{2C\|\mathbf{w}\|_{L^{q}}}\right)^{1/(\beta+\alpha+\sigma)}\right\},$
we find that
$t^{\beta}\|\mathcal{L}_{u_{0}}(t,\cdot)\|_{L^{q}}\leq R_{0}^{\beta}\|u_{0}\|_{L^{q}}\leq\frac{\mathscr{M}}{2},\quad\text{and}\quad t^{\beta}\|\mathcal{S}_{\mathbf{w}}(t,\cdot)\|_{L^{q}}\leq CR_{0}^{\beta+\alpha+\sigma}\|\mathbf{w}\|_{L^{q}}\leq\frac{\mathscr{M}}{2}.$
This establishes the desired bound on the interval $0,R_{0}]$. For $t>R_{0}$, we analyze the behavior by decomposing the spatial domain into three regions:
$\mathcal{E}_{1}$ $=\{y\in\mathbb{R}^{N}:|y|>Mt^{\frac{\alpha}{2\mathsf{s}}}\},$
$\mathcal{E}_{2}$ $=\{y\in\mathbb{R}^{N}:|y|<Mt^{\frac{\alpha}{2\mathsf{s}}},\ |x-y|<\delta t^{\frac{\alpha}{2\mathsf{s}}}\},$
$\mathcal{E}_{3}$ $=\{y\in\mathbb{R}^{N}:|y|<Mt^{\frac{\alpha}{2\mathsf{s}}},\ |x-y|>\delta t^{\frac{\alpha}{2\mathsf{s}}}\}.$
We decompose the linear part as $\mathcal{L}_{u_{0}}=\mathcal{L}_{u_{0},1}+\mathcal{L}_{u_{0},2}+\mathcal{L}_{u_{0},3}$, where each term corresponds to integration over $\mathcal{E}_{1}$, $\mathcal{E}_{2}$, $\mathcal{E}_{3}$ respectively. Similarly, we write $\mathcal{S}_{\mathbf{\mathbf{w}}}=\mathcal{S}_{\mathbf{\mathbf{w},1}}+\mathcal{S}_{\mathbf{\mathbf{w},2}}+\mathcal{S}_{\mathbf{\mathbf{w},3}}$, with
$\mathcal{S}_{\mathbf{\mathbf{w},j}}(t,x)=\int\limits_{0}^{t}\tau^{\sigma}\int\limits_{\mathcal{E}_{j}}\mathbf{\mathbf{w}}(y)Y(t-\tau,x-y)dyd\tau,\quad j={1,2,3}.$
We now estimate each term separately. For $\mathcal{L}_{u_{0},1}$, using (1.6) and the kernel estimate $\|Z(t,\cdot)\|_{L^{r}}=\|F\|_{L^{r}}t^{-\frac{N\alpha}{2\mathsf{s}}(1-\frac{1}{r})}$, we obtain
$t^{\beta}\|\mathcal{L}_{u_{0},1}(t,\cdot)\|_{L^{q}}\leq t^{\beta}\|u_{0}\|_{L^{p_{c}}(\mathcal{E}_{1})}\|Z(t,\cdot)\|_{L^{r}}\leq\frac{\mathscr{M}}{6}.$
For $\mathcal{S}_{\mathbf{w},1}$, condition (1.7) and the bound $\|Y(t,\cdot)\|_{L^{r}}\leq Ct^{1-\alpha-\frac{N\alpha}{2\mathsf{s}}(1-\frac{1}{r})}$ give
$t^{\beta}\|\mathcal{S}_{\mathbf{w},1}(t,\cdot)\|_{L^{q}}\leq t^{\beta}\|\mathbf{w}\|_{L^{p_{c}}(\mathcal{E}_{1})}\int\limits_{0}^{t}\tau^{\sigma}\|Y(t-\tau,\cdot)\|_{L^{r}}d\tau\leq\frac{\mathscr{M}}{6}.$
The local terms $\mathcal{L}_{u_{0},2}$ and $\mathcal{S}_{\mathbf{\mathbf{w},2}}$ are controlled via (1.8) and the local boundedness of $F$ and $G$:
$t^{\beta}\|\mathcal{L}_{u_{0},2}(t,\cdot)\|_{L^{q}}\leq\frac{\mathscr{M}}{6},\quad t^{\beta}\|\mathcal{S}_{\mathbf{\mathbf{w},2}}(\cdot,t)\|_{L^{q}}\leq\frac{\mathscr{M}}{6}.$
Finally, the nonlocal terms $\mathcal{L}_{u_{0},3}$ and $\mathcal{S}_{\mathbf{\mathbf{w},3}}$ are estimated using (1.9) and (1.10), which exploit the tail behavior of $F$ and the averaged smallness of $u_{0}$ and $\mathbf{\mathbf{w}}$:
$t^{\beta}\|\mathcal{L}_{u_{0},3}(t,\cdot)\|_{L^{q}}\leq\frac{\mathscr{M}}{6},\quad t^{\beta}\|\mathcal{S}_{\mathbf{\mathbf{w},3}}(t,\cdot)\|_{L^{q}}\leq\frac{\mathscr{M}}{6}.$
Summing these bounds yields
$t^{\beta}\left(\|\mathcal{L}_{u_{0}}\|_{L^{q}}+\|\mathcal{S}_{\mathbf{w}}\|_{L^{q}}\right)\leq\mathscr{M}\quad\text{for }t>R_{0}.$
The global existence of the solution follows from Lemma 2.6, which concludes the proof.
Declarations. On behalf of all authors, the corresponding author states that there is no conflict of interest. No data-sets were generated or analyzed during the current study.